OpenStudy (anonymous):

(4+√50)-(3-√8) a. 1-3√2 b.1+7√2 c. 1+√58 d. 1+√42

OpenStudy (anonymous):

Its B.

OpenStudy (helder_edwin):

u must know this \[ \large \sqrt{50}=\sqrt{25\cdot2}=\sqrt{25}\sqrt{2}=5\sqrt{2} \]

OpenStudy (anonymous):

Yes Thanks @helder_edwin

do the same with \(\sqrt{8}\):